# Grounding Multimodal LLMs to Embodied Agents that Ask for Help with Reinforcement Learning

### v2

@@ 4 Approach
Figure 2: MLLM Policy Architecture. The policy takes as input a task instruction, sequence of past observations, actions, user response to questions asked and outputs a high-level action (i.e. skill) or a question in natural language.
Our method adapts a pre-trained multi-modal large language model (MLLM) to a vision-language-action (VLA) model capable of simultaneously taking actions, and generating natural language questions to resolve ambiguity in multi-modal embodied decision-making settings – without relying on human annotations or hand-engineered rewards. We use reinforcement learning (RL) with LLM-generated rewards to train a VLA for Ask-to-Act.
@@ 4.1 Policy Architecture
We adapt the LLaVA-OneVision [22] MLLM architecture to a VLA model for task-planning and asking questions as illustrated in Fig. 2. The MLLM policy takes as input a task instruction along with a sequence of past observations, actions, user response to questions asked and outputs a high-level action (i.e. skill) or a question in natural language. As shown in Fig. 2, the task instruction is first encoded into language tokens. Traditional MLLMs embed a single visual input into a large number of visual tokens e.g. LLaVA-OneVision [22] uses $729$ tokens. However, in partially observable embodied settings these agents need to reason over a long sequence of past observations to complete a task. Using a large number of visual tokens for each observation would quickly increase the context lengths (e.g. $32$ visual observations ×$\times 729$ tokens $={\sim}23k$ tokens). To make it feasible to train MLLMs with a long observation history, we employ a Perceiver model [19] to downsample the visual tokens per observation, similar to [41]. Specifically, we encode each visual observation using the MLLM visual encoder and then pass output to a Perceiver downsampler which outputs $k<729$ tokens ($k=4$ in our case). Next, these downsampled visual tokens from the current observation $o_{t}$ are interleaved with task instruction tokens ($l1,...,l_{k}$), text representation of actions token ($u^{1}_{t-1},..,u^{m}_{t-1}$), and language response given by user for any questions asked at each step tokenized ($h^{1}_{t-1},..,h^{k}_{t-1}$).
These interleaved language and visual tokens are processed by the MLLM, which autoregressively predicts a sequence of language tokens corresponding to a discrete action or a natural language question. Each agent action $a\in A$ (a skill or possible questions $q\in Q$) is represented as free-form text sequence, tokenized as $(u^{1}_{t},...,u^{m}_{t})$. For instance, in an action space where the high-level skill is described as “pick red bowl”, the sequence is tokenized as $[29245,2518,19212]$ using the Qwen-2 [51] tokenizer. The MLLM must predict tokens $29245$, then $2518$, and finally $19212$ to execute the action. Sequences corresponding to invalid actions are either filtered out or treated as a no-op to prevent execution errors.
Constrained Grammar Decoding. Training MLLMs with RL using language actions significantly expands the action space where most actions are invalid, which makes RL training difficult [44]. For instance, the Qwen-2 [51] tokenizer has a vocabulary of $151646$ tokens, resulting in $151646^{m}$ possible action predictions per step, where only a small subset of these actions are valid. To address this, we employ grammar-constrained decoding [29] to limit autoregressive sampling to only valid action token sequences. Grammar-Constrained Decoding is defined as a function $M(u^{1}_{t},...,u^{j-1}_{t})$ that generates a binary mask over all tokens, indicating which tokens are valid for the $j$th decoding step. This restriction ensures that model only predicts valid actions, making RL training tractable.
@@ 4.2 Training with Reinforcement Learning using LLM-generated Rewards
To finetune LLaVA-OneVision [22] base MLLM with embodied data we use online RL with LLM-generated rewards for Ask-to-Act task. Manually defining dense rewards using heuristics to generate subgoals and evaluate questions within the task context is not trivial. We argue that evaluating natural language questions and agent actions for such tasks requires writing programs that can evaluate task-specific deductive and contextual reasoning, as well as language generation, which is not feasible. We show that current state-of-the-art LLMs, when provided with task and environment information in the right representation, demonstrate contextual understanding and deductive reasoning ability required to reason about such ambiguity in embodied settings. This suggests that LLMs can be used to automatically generate subgoals, and evaluate whether a sequence of questions can disambiguate the target object for Ask-to-Act, which can be used as reward signal for RL training. Specifically, we prompt an LLM with privileged environment state, task instruction, and target object metadata in a text representation, and ask it to generate relevant subgoals and all optimal question sequences to solve each Ask-to-Act task. See Appendix D for more details and the prompt.
We use generated subgoals and question sequences to generate the reward at each timestep using the equation described in Eq. 1. The reward function consists of a sparse reward for completing the task, subgoal rewards for achieving intermediate objectives, a slack penalty to encourage task completion in fewer steps, and a reward for asking clarification questions that aids in disambiguating the target object. The reward at step $t$ is computed as follows:
@@
where $\mathds{1}_{\text{success}}$ indicates whether the task was successfully completed and $\mathds{1}_{\text{subgoal}}$ (generated by the LLM) indicates if the agent completed any subgoal required to complete the overall task. For example, for a task “Bring me the cup and place it on the coffee table”, the agent needs to first search for all cups, pick the correct cup, then navigate to the coffee table, and finally place it. Similarly, $\mathds{1}_{\text{useful\_question}}$ (generated by the LLM) indicates if a question asked by the agent is valid and helps make progress towards disambiguating the target object or not. Consider the example of fetching the cup, if the environment has $4$ cups on a table and the agent asks “Is it the red cup?” and if user responds ‘no’, then the agent should ask ‘Is it the blue cup?’ or ‘Is it the yellow cup?’ to find the target object instead of asking ‘Is it on the table?’ to make progress towards solving the task. Additionally, for each question the agent asks in an episode the reward $r_{3}$ is a value normalized to sum up to $1$ based on number of questions required to solve the task. When training policies under a budget of questions we penalize the agent for every question that exceeds pre-specified budget given by $\mathds{1}_{\text{exceed\_budget}}$. By default the budget is set to minimum number of required questions $K$ to solve the task.
Implementation details. We train the MLLM policy using DD-PPO [47], an adaptation of PPO [36] for distributed training, for $50$ million steps on $8$ A40 GPUs. To generate rewards for our task we use the Llama-3 [16] LLM. We provide full training details in Appendix C.
@@ 5 Experiments
In this section, we compare our method of adapting MLLMs as VLA using RL with LLM-generated rewards to proprietary MLLMs in a zero-shot setting and MLLMs trained using synthetically generated SFT data for the our task.
@@ 5.1 Baselines
For methods involving MLLM fine-tuning, we use LLaVA-OneVision $0.5$[22] as our base MLLM with modifications described in Sec. 4.1. We use the $0.5$B model due to compute constraints for running RL training on A40 GPUs. Below, we outline the details of each baseline:
@@
(a.) Fully Observable Text WorldGraph + ReAct (Zero-shot). In this baseline, we provide an LLM (GPT4o in our case) with a fully observable text-based world graph describing the environment, including receptacles, objects, and their locations (e.g. “The apple is on the coffee table.”). Additionally, the LLM is equipped with an skill library for executing actions, a history of previous actions, and the task instruction. At each timestep, the LLM generates a reasoning chain via ReAct [53] prompting followed by an action to complete the task. By eliminating perception limitations, this baseline evaluates the zero-shot planning and reasoning capabilities of LLMs for solving ambiguous embodied tasks.
@@
(b.) Fully Observable Text WorldGraph + ReAct (Few-shot). This baseline extends (a) by providing the LLM with a few in-context examples that demonstrate task planning and ambiguity resolution strategies in the Ask-to-Act task. By leveraging demonstrations, this approach assesses whether in-context learning improves LLM’s task planning and ambiguity reasoning under full observability.
@@
(c.) Partially Observable Text WorldGraph + ReAct (Few-shot). While baselines (a) and (b) assume privileged access to a fully observable world graph, constructing such representations in real-world settings is often infeasible. This baseline relaxes that assumption by providing the LLM with a partially observable text-based world graph. At the start of an episode, the agent lacks full knowledge of object locations and must actively explore to gather necessary information.
@@
(d.) Vision GPT4o + SoM + ReAct. Building an error-free text representation of real-world environments is challenging. This baseline evaluates whether existing MLLMs can solve the Ask-to-Act task using egocentric visual observations. At each timestep, the MLLM receives the robot’s visual input along with an skill library for executing actions. To enhance grounding, we label visual observations using Set-of-Marks (SoM) [52] and maintain memory by providing GPT4o with a textual history of past observations and actions.
@@
(e.) LLaVA-OneVision SFT. This baseline finetunes LLaVA-OneVision $0.5$B MLLM for our task using Supervised Fine-Tuning (SFT) on synthetically generated trajectories. Generating training data is non-trivial, as the agent must first explore the environment to locate relevant objects, ask contextually grounded questions to disambiguate the target object, and finally execute the instruction once the object is identified. To achieve this, we design an expert using heuristic exploration and an LLM. The agent first explores the scene using frontier exploration [50] until all receptacles are seen, then converts its observations into a text-based world graph. This graph, along with the task instruction and privileged information about the target object, is passed to an LLM, which generates a sequence of clarifying questions to identify the object. Once the target object is determined, a heuristic planner executes the rearrangement task. The final trajectory consists of actions taken during exploration, LLM-generated questions as actions, and actions taken to rearrange the object. Only successful trajectories are included in training. Our SFT dataset contains ${\sim}40$k trajectories in $63$ training scenes in the ReplicaCAD dataset.
@@ 5.3 Results
[…]
1) WG + ReAct* $56.1$ $50.5$ $3.2$ $47.3$ $37.4$ $2.3$
2) WG + ReAct* (FS) $96.2$ $49.7$ $2.3$ $94.7$ $35.6$ $1.9$
3) WG + ReAct* (FS) $61.8$ $46.9$ $3.2$ $49.2$ $32.2$ $2.2$
4) GPT4o + SoM + ReAct $25.2$ $20.1$ $1.9$ $19.8$ $12.5$ $1.1$
5) LLaVA-OV SFT $48.2$ $46.9$ $\mathbf{1.5}$ $34.1$ $26.1$ 0.8
6) LLaVA-OV RL (Ours) $\mathbf{89.8}$ $\mathbf{63.2}$ $2.6$ $\mathbf{65.2}$ $\mathbf{32.4}$ $2.5$
@@
Table 2: Results. Evaluation of all methods on Unseen Scenes and Unseen Tasks evaluation splits of Ask-to-Act task. FS denotes few-shot examples, * denotes access to privileged information, Full Obs. stands for full observability.
@@
Tab. 2 presents the results of evaluating all methods on Unseen Scenes and Unseen Tasks splits. To establish an upper bound on performance, we first evaluate zero-shot LLMs with privileged information (rows 1 and 2 of Tab. 2). When LLMs are provided with a fully observable text-based representation of the environment and use ReAct [53] for action prediction, they achieve a success rate of $56.1\%$ and $47.3\%$ and an ARS of $50.5\%$ and $37.4\%$ on Unseen Scenes and Unseen Tasks, respectively. This highlights the inherent difficulty of task planning and ambiguity resolution for LLMs in a zero-shot setting. Next, we augment LLMs with privileged in-context examples demonstrating how to perform the Ask-to-Act task. This significantly improves performance, increasing success rate to $96.2\%$ and $94.7\%$ on unseen scenes and tasks. These promising results motivated us to explore using LLMs as reward generators for Ask-to-Act. We then evaluate the same baseline under partial observability (row 3), where the agent does not have access to a fully observable world graph but still has in-context examples. This leads to an absolute drop in success rate of $34.4\%$ and $45.5\%$ on both splits (row 3 vs. row 2), indicating that task planning and ambiguity resolution become significantly more challenging when the agent must actively explore to gather information. As shown in Tab. 2, our approach (row 6), which fine-tunes a MLLM [22] using RL with LLM-generated rewards, outperforms all methods operating under partial observability and achieves $89.8\%$ and $65.2\%$ on success rate and $63.2\%$ and $32.2\%$ ARS on Unseen Scenes and Unseen Tasks splits. Our method outperforms the policy trained with synthetic SFT data (row 5 vs. row 6) by absolute $31.1-41.6\%$ on success rate and $6.3-16.3\%$ on ARS. We find agents trained with SFT tend to be more conservative (lower than $1$ QR as shown in row 5), asking fewer questions compared to RL-trained counterparts which could be the reason behind difference in performance, additional analysis shown in Sec. 5.4. Additionally, our method surpasses a strong zero-shot baseline (row 4) that uses vision GPT4o [26] with Set-of-Marks (SoM) [52] prompting and a text-based history of past observations by an absolute margin of $45.4-64.6\%$ on success rate. These results demonstrate online RL combined with LLM-generated rewards is an effective method for training embodied agents that can interleave acting with asking relevant clarification questions to resolve ambiguity.
Can agents succeed without asking for help? In this experiment, we ask whether it is possible to train an effective MLLM policy using RL on Ask-to-Act task without asking a single question. As shown by results in Tab. 5 in Appendix B, we find agents trained to solve our task without asking questions achieve performance close to a random baseline (row 1 vs. row 2). In contrast, a policy trained using our LLM generated rewards with RL achieves significantly better performance on both Unseen Scenes and Unseen Tasks evaluation splits. More details in Appendix B.
@@
[…]
@@
8) Subgoal Reward $34.9$ $30.7$ $3.8$ $16.5$ $6.9$ $2.9$
9) Ours $\mathbf{89.8}$ $\mathbf{63.2}$ $\mathbf{2.6}$ $\mathbf{65.2}$ $\mathbf{32.4}$ $\mathbf{2.5}$
@@
Choice of reward for RL training. In Tab. 3, we ablate the choice of reward function for RL training to demonstrate effectiveness of using a LLM to generate rewards for Ask-to-Act. We consider two simple rewards that can be programmatically defined for our task:
1.) Success reward: a simple sparse reward setting where the agent is only given a reward of value $10$ at the end of the episode if the agent succeeds and 00 otherwise. In addition, the agent also receives a slack penalty of $0.01$ at each timestep to encourage faster task execution.
@@
2.) Subgoal reward: in this setting the agent is rewarded for accomplishing any subgoal that is required to completing the overall task. For example, for a task “Bring me the cup and place it on the coffee table”, the agent should navigate to the target cup, pick the target cup, then navigate to coffee table, and finally place it. In this case, the agent will be rewarded for navigating to only the target object instance, picking the target object, navigating to target receptacle and placing it. This reward can be programmatically generated for training episodes with access to privileged information from simulator and target object metadata from each episode. Note, this reward does not provide dense step-by-step incentive for exploration or to ask a sequence of questions required to disambiguate the object; instead, it implicitly rewards the agent to ask questions that lead to the agent picking the right object which makes RL training using reward difficult.
As shown in Tab. 3, neither of these rewards are sufficient to learn an effective MLLM policy for our task, demonstrating the need for precise per-step rewards that incentivize agents to not only act but also ask meaningful questions.
@@
Figure 3: Task Performance vs. Budget of Questions. Evaluation performance of policies trained under different budget of questions vs. task Success Rate and Ambiguity Resolution Efficiency score.
@@
Training policies under variable budget. A desired skill for an embodied agent that can ask questions is to adhere to user’s preferences about how often they would like a robot to ask clarification questions. Some would prefer an agent ask as few questions as possible for better user experience by trading off task success rates. In contrast, some users would be fine with an agent asking as many questions as it would like to ensure task success rates are higher. Motivated by this, we train multiple MLLM policies using RL with LLM generated rewards with a variable upper-bound on maximum number of questions an agent can ask. Specifically, we train policies with a budget of $B\in\{K,K+1,K+2,K+4\}$ questions, where $K$ is defined as minimum required question for a task in Ask-to-Act dataset. In this setting, an agent can ask at most $B$ questions in a single episode (either relevant or irrelevant) without incurring any penalties. Note, for this experiment $B$ is either equal to $K$ i.e. ask as close to minimum required questions as possible or can be quite high $K+4$ where an agent can ask as many as $4$ extra questions than minimum required in each episode without incurring any penalties. Additionally, the agent will only be rewarded for relevant questions from all questions it asked. We only penalize the agent for each question asked after exceeding the question budget $B$.  Fig. 3 shows success rates of various policies trained with different budgets under the reward setting described in Eq. 1. As shown in Fig. 3 (a.), as we increase the number of questions the agent can ask, the success rates increase; however, there is a clear trade-off between increase in success rates and question ratio (i.e. asked questions to minimum required), see  Fig. 3 (b.).

### v3

@@ 4 Approach
Figure 2: MLLM Policy Architecture. The policy takes as input a task instruction, past observations, actions, user response to questions asked and outputs a high-level action or a question in natural language.
Our method, called AutoAsk, adapts a pre-trained multi-modal large language model (MLLM) to a vision-language-action (VLA) model capable of simultaneously taking actions, and generating natural language questions to resolve ambiguity in multi-modal embodied decision-making settings – without relying on human annotations or hand-engineered rewards. We use reinforcement learning (RL) with LLM-generated rewards to train a VLA for Ask-to-Act.
@@ 4.1 Policy Architecture
We adapt the LLaVA-OneVision (Li et al., 2024) MLLM architecture to a VLA model for task-planning and asking questions as illustrated in Fig.˜2. The MLLM policy takes as input a task instruction along with a sequence of past observations, actions, user response to questions asked and outputs a high-level action (i.e. skill) or a question in natural language. As shown in Fig.˜2, the task instruction is first encoded into language tokens. Traditional MLLMs embed a single visual input into a large number of visual tokens e.g. LLaVA-OneVision (Li et al., 2024) uses $729$ tokens. However, in partially observable embodied settings these agents need to reason over a long sequence of past observations to complete a task. Using a large number of visual tokens for each observation would quickly increase the context lengths (e.g. $32$ visual observations $\times 729$ tokens ${\sim}23k$ tokens). To make it feasible to train MLLMs with a long observation history, we employ a Perceiver model (Jaegle et al., 2022) to downsample the visual tokens per observation (to $k=4$ in our case), similar to Szot et al. (2024a). Next, these downsampled visual tokens from the current observation $o_{t}$ are interleaved with task instruction tokens ($l_{1},...,l_{k}$), text representation of past actions ($u^{1}_{t-1},..,u^{m}_{t-1}$), and language response given by user for questions asked at each step tokenized ($h^{1}_{t-1},..,h^{n}_{t-1}$). These interleaved language and visual tokens are processed by the MLLM, which autoregressively predicts a sequence of language tokens corresponding to a discrete action or a natural language question. Each agent action $a\in A$ (a skill or possible questions $q\in Q$) is represented as free-form text sequence, tokenized as $(u^{1}_{t},...,u^{m}_{t})$. Sequences corresponding to invalid actions are treated as a no-op to prevent execution errors.
@@ 4.2 Training with Reinforcement Learning using LLM-generated Rewards
To finetune LLaVA-OneVision (Li et al., 2024) base MLLM with embodied data we use online RL with LLM-generated rewards for the Ask-to-Act task. We show that current state-of-the-art LLMs can be used to generate reward signal for RL training, when provided with task and environment information in the right representation. Specifically, when the agent asks a question we provide an LLM with privileged environment state, task instruction, metadata of all target objects, expected target location, questions asked so far and action at current step in text representation. Next, we ask the LLM to generate a binary reward that evaluates whether the question asked helps resolve ambiguity or identify user preference compared to information from the previous step. See Appendix˜C for more details and the prompt. Eq.˜1 describes the reward function we use, which consists of a sparse reward $\mathds{1}_{\text{success}}$ for completing the task, subgoal rewards $\mathds{1}_{\text{subgoal}}$ generated by the LLM for achieving intermediate objectives (e.g. picking target object, navigating to target location, etc), a ckck penalty $r_{5}$ at every timestep to encourage task completion in fewer steps, and a reward for asking clarification questions $\mathds{1}_{\text{useful\_question}}$ that aids in disambiguating the target object. The reward at step $t$ is computed as follows:
@@
where $\mathds{1}_{\text{useful\_question}}$ (computed using an LLM as a reward model) indicates if a question asked by the agent is valid and helps make progress towards disambiguating the target object or not. Consider the example of placing utensils in kitchen, if the environment has $2$ cups and $1$ plates on a table and the agent asks “Where should I place cups?” and if the user responds “left drawer of kitchen counter”, then the agent should ask “Where should I move the plates?” to make progress towards the task instead of asking “Is the plate on the table?”. When training policies under a budget of questions we penalize the agent for every question that exceeds a pre-specified budget given by $\mathds{1}_{\text{exceed\_budget}}$. By default the budget is set to the minimum number of required questions $K$ to solve the task.
Implementation details. We train the MLLM policy using DD-PPO (Wijmans et al., 2020), an adaptation of PPO (Schulman et al., 2017) for distributed training, for $50$ million steps on $8$ A40 GPUs. To generate rewards for our task we use the Llama-3 (Grattafiori et al., 2024) LLM. The LLM reward model runs as a separate process using vLLM (Kwon et al., 2023) and processes reward labeling requests asynchronously. We provide training details and prompts in Appendix˜B and Appendix˜F.
@@ 5 Experiments
In this section, we compare our method (AutoAsk) of adapting MLLMs as VLA using RL with LLM-generated rewards to proprietary MLLMs in a zero-shot setting and MLLMs trained using synthetically generated SFT data for our task.
@@ 5.1 Baselines
For methods involving MLLM fine-tuning, we use LLaVA-OneVision 0.50.5B (Li et al., 2024) as our base MLLM with modifications described in Sec.˜4.1. We use the 0.50.5B model due to compute constraints for running RL training on A40 GPUs. Below, we outline the details of each baseline:
@@
(a.) Fully Observable Text WorldGraph + ReAct (Zero-shot). We provide GPT4o with a fully observable text-based world graph describing the environment, including receptacles, objects, and their locations (e.g. “The apple is on the coffee table.”). Additionally, the LLM is equipped with a skill library for executing actions, a history of previous actions, and the task instruction. At each timestep, the LLM generates a reasoning chain via ReAct (Yao et al., 2023) prompting followed by an action to complete the task. By eliminating perception limitations, this baseline evaluates the zero-shot planning and reasoning capabilities of LLMs for solving under-specified embodied tasks.
@@
(b.) Fully Observable Text WorldGraph + ReAct (Few-shot). This baseline extends (a) by providing the LLM with a few in-context examples that demonstrate task planning and ambiguity resolution strategies in Ask-to-Act to assesses whether in-context learning improves task performance.
@@
(c.) Partially Observable Text WorldGraph + ReAct (Few-shot). While baselines (a) and (b) assume privileged access to a fully observable world graph, constructing such representations in real-world settings is often infeasible. This baseline relaxes that assumption by providing the LLM with a partially observable text-based world graph requiring LLM agent to actively explore to solve the task.
@@
(d.) Vision GPT4o + SoM + ReAct. Building an error-free text representation of real-world environments is challenging. This baseline evaluates whether existing MLLMs can solve the Ask-to-Act task using egocentric visual observations. At each timestep, MLLM receives the robot’s visual input annotated with Set-of-Marks (SoM) (Yang et al., 2023) along with a skill library and a memory module that maintains textual history of past observations and actions.
@@
(e.) LLaVA-OneVision SFT. Generating training data for SFT is non-trivial, as the agent must first explore the environment to locate relevant objects, ask contextually grounded questions at the right time to resolve ambiguity and identify user preferences, and finally execute the task. To achieve this, we design an expert using heuristic exploration and an LLM. The agent first explores the scene using frontier exploration (Yamauchi, 1997) until all receptacles are seen, then converts its observations into a text-based world graph. This graph, along with the task instruction and privileged information about the target objects, is passed to an LLM, which generates a sequence of clarifying questions to identify target objects. Next, a heuristic planner executes the rearrangement tasks to generate the final trajectory which consists of exploration actions, LLM-generated questions as actions, and actions taken to rearrange objects. Only successful trajectories are included in training. Our SFT dataset contains ${\sim}40$k trajectories in $63$ training scenes in the ReplicaCAD dataset.
@@ 5.3 Results
Method Obs. SR $(\mathbf{\uparrow})$ ARS $(\mathbf{\uparrow})$ QR $(\mathbf{\downarrow})$ SR $(\mathbf{\uparrow})$ ARS $(\mathbf{\uparrow})$ QR $(\mathbf{\downarrow})$
1) WG + ReAct* $41.7$ $38.9$ $3.3$ $39.9$ $37.1$ $3.2$
2) WG + ReAct* (FS) $46.7$ $44.7$ $0.7$ $44.4$ $42.6$ $1.2$
3) WG + ReAct* (FS) $35.5$ $26.6$ $3.9$ $32.2$ $26.4$ $3.8$
4) GPT4o + SoM + ReAct $15.7$ $13.5$ $1.9$ $14.3$ $12.7$ $2.1$
5) LLaVA-OV SFT $33.6$ $26.6$ $\mathbf{0.8}$ $29.5$ $25.1$ $\mathbf{1.1}$
6) AutoAsk (Ours) $\mathbf{45.9}$ $\mathbf{40.6}$ $0.9$ $\mathbf{38.7}$ $\mathbf{35.2}$ $1.5$
@@
Table 1: Results. Evaluation of all methods on Unseen Scenes and Unseen Tasks evaluation splits. FS denotes few-shot examples, * denotes access to privileged information, Full Obs. stands for full observability.
@@
Tab.˜1 presents the results of evaluating all methods on Unseen Scenes and Unseen Tasks splits. To establish an upper bound on performance, we first evaluate zero-shot LLMs with privileged information (rows 1 and 2 of Tab.˜1). When an LLM is provided with a fully observable text-based representation of the environment and uses ReAct (Yao et al., 2023), it achieves a success rate of $41.7\%$ and $39.9\%$ and ARS of $38.9\%$ and $37.1\%$ on Unseen Scenes and Unseen Tasks splits. This highlights the inherent difficulty of task planning and ambiguity resolution for LLMs in a zero-shot setting. Next, we augment LLMs with privileged in-context examples demonstrating how to perform the Ask-to-Act tasks. This significantly improves performance, increasing success rate to $46.7\%$ and $44.4\%$ on unseen scenes and tasks. With in-context examples we find LLMs perform quite well on single-object rearrangement tasks but struggle with multi-object rearrangement tasks where interaction with articulated objects is required. We then evaluate the same baseline under partial observability (row 3), where the agent does not have access to a fully observable world graph but still has in-context examples. This leads to an absolute drop in success rate by $11.2\%$ and $12.2\%$ on both splits (row 3 vs. row 2), indicating that task planning and ambiguity resolution become significantly more challenging when the agent must actively explore to gather information. As shown in Tab.˜1, our approach AutoAsk (row 6), which fine-tunes a MLLM (Li et al., 2024) using RL with LLM-generated rewards, outperforms all methods operating under partial observability and achieves $45.9\%$ and $38.7\%$ on success rate and $40.6\%$ and $35.2\%$ ARS on Unseen Scenes and Unseen Tasks splits. AutoAsk outperforms the policy trained with synthetic SFT data (row 5 vs. row 6) by absolute $6.5-12.3\%$ on success rate and $8.8-20.0\%$ on ARS. We find agents trained with SFT tend to be more conservative (lower than $1$ QR as shown in row 5), asking fewer questions compared to RL-trained counterparts which could be the reason for difference in performance, additional analysis shown in Sec.˜5.4. Additionally, AutoAsk also surpasses a strong zero-shot baseline (row 4) that uses vision GPT4o (OpenAI, 2023) with SoM (Yang et al., 2023) prompting and a text-based history of past observations by an absolute margin of $24.4-30.2\%$ on success rate. These results demonstrate online RL combined with LLM-generated rewards is an effective method for training agents that can interleave acting with asking relevant clarification questions to resolve ambiguity.
@@
Method SR $(\mathbf{\uparrow})$ ARS $(\mathbf{\uparrow})$ QR $(\mathbf{\downarrow})$ SR $(\mathbf{\uparrow})$ ARS $(\mathbf{\uparrow})$ QR $(\mathbf{\downarrow})$
@@
8) Subgoal Reward $22.4$ $21.6$ $\mathbf{0.4}$ $16.5$ $6.9$ $\mathbf{0.5}$
9) AutoAsk $\mathbf{45.9}$ $\mathbf{40.6}$ $0.9$ $\mathbf{38.7}$ $\mathbf{35.2}$ $1.5$
@@
Choice of reward for RL training. In Tab.˜2, we ablate the choice of reward function for RL training to demonstrate effectiveness of using an LLM to generate rewards for Ask-to-Act. We consider two simple rewards that can be programmatically defined for our task: (1.) Success reward: a sparse success reward of value $10$ only given a at the end of the episode if the agent succeeds and 0 otherwise. (2.) Subgoal reward: the agent is rewarded for accomplishing any subgoal $\mathds{1}_{\text{subgoal}}$ that are required to completing the overall task as defined in Eq.˜1 in addition to a success reward. This reward can be programmatically generated for training episodes with access to privileged information from simulator for each episode. Note, this reward does not provide dense step-by-step incentive for exploration or to ask questions to resolve ambiguity; instead, it implicitly rewards the agent to ask questions that lead to subgoal success which makes RL training difficult. As shown in Tab.˜2, neither of these rewards are sufficient to learn an effective MLLM policy for our task, demonstrating the need for precise per-step rewards that incentivize agents to not only act but also ask meaningful questions.
@@
Figure 3: Analysis. (a.) Task Performance vs. Budget of Questions. Evaluation performance of policies trained under different budget of questions vs. task Success Rate and Ambiguity Resolution Efficiency score. (c.) Success Rate vs. Number of Required and Asked Questions. Evaluation performance of SFT and RL trained policies vs. number of required and total questions asked by the agent.
@@
Training policies under variable budget. A desired skill for an embodied agent that can ask questions is to adhere to user’s preferences about how often they would like a robot to ask clarification questions. Some would prefer an agent ask as few questions as possible for better user experience by trading off task success rates. In contrast, some users would be fine with an agent asking as many questions as it would like to ensure task success rates are higher. Motivated by this, we train multiple MLLM policies using RL with LLM generated rewards with a variable upper-bound on maximum number of questions an agent can ask. Specifically, we train policies with a budget of $B\in\{K,K+1,K+2,K+4\}$ questions, where $K$ is minimum required question for a task in Ask-to-Act dataset. The agent can ask at most $B$ questions in a single episode (either relevant or irrelevant) without incurring any penalties. Note, for this experiment $B$ is either equal to $K$ i.e. ask as close to minimum required questions as possible or can be quite high $K+4$ where an agent can ask as many as $4$ extra questions than minimum required in each episode without incurring any penalties. Additionally, the agent will only be rewarded for relevant questions from all questions it asked. We only penalize the agent for each question asked after exceeding the question budget $B$. Fig.˜3 (a.) and (b.) shows success rates and question ratio of policies trained with different budgets under the reward setting described in Eq.˜1. As shown in Fig.˜3 (a.), as we increase the number of questions the agent can ask, the success rates increase; however, there is a clear trade-off between increase in success rates and question ratio (i.e. asked questions to minimum required), see Fig.˜3 (b.).